# Ensuring Force Safety in Vision-Guided Robotic Manipulation via Implicit Tactile Calibration

### v1

@@ IV Dataset
V Benchmark
V-A Evaluation Metrics
@@
Success Rate (SuR) Following [18], we leverage SuR to quantify the effectiveness of the state planning model in robotic manipulation tasks, by calculating the proportion of test scenarios in which the model can successfully complete the task out of the total number of test scenarios.
Safety Rate (SaR-95) and Sub-Safety Rate (SaR-80) Safety Rates are utilized to evaluate the scenario-wise force safety of the state planning model in robotic manipulation tasks. A test scenario is considered safely manipulated only if the harmful forces magnitude $\|\textbf{F}_{\texttt{harmful}}\|$ within the whole continuous manipulation trajectory remain below a specific force magnitude threshold f at all times. But it is worth mentioning that the force safety of a state planning model only depends on the state generated by itself, rather than other states in that manipulation trajectory. Thus, we discretize the above defined as,
$\|\textbf{F}_{\texttt{harmful}}\|^{k}<\textbf{f},\quad\forall k\in[1,L]$ (6)
where $L$ denotes the length of states planned by the model. To eliminate the impact of noise and make the safety evaluation more robust, we relaxed the above definition to the following two metrics: relaxed safety and sub-safety. Concretely, a test scenario can be considered manipulated by the model in a relaxed safe manner when 95% of its generated states have the harmful force below f. Therefore, the safety of the state planning model can be computed as,
[…]
where $\textbf{Num}_{\texttt{safe}}$ denotes the number of test scenarios which is manipulated successfully in a relaxed safe manner, and $\textbf{Num}_{\texttt{success}}$ denotes the total number of test scenarios which is manipulated successfully. Similarly, a test scenario can be considered manipulated by the model in a sub-safe manner when 80% of its generated states have the harmful force below f. The sub-safety of the state planning model can be computed as,
[…]
where $\textbf{Num}_{\texttt{sub-safe}}$ denotes the number of test scenarios which is manipulated successfully in a sub-safe manner.
@@
Average Harmful Force (AHF) AHF is applied to evaluate the force-wise force safety of the state planning model in robotic manipulation tasks. It is calculated as the average harmful force magnitude $\|\textbf{F}_{\texttt{harmful}}\|$ applied throughout the robot manipulation process across all test scenarios.
V-B Simulation Experiments
@@
Implementation Our proposed SafeDiff model is implemented based on the publicly available Decision Diffuser code base [1]. The training and testing processes are conducted using an NVIDIA A100 Tensor Core GPU. We utilize the training demonstrations provided by our SafeDoorManip50k for safe state planning. The training configuration is as follows: batch size is $256$, total training epochs are $500$, an initial learning rate of $10^{-4}$ with a decay rate of $0.985$, and the application of an Exponential Moving Average (EMA) with a decay factor of $0.995$. During testing, we evaluate the performance of the safe state planning models under $4,580$ seen-door scenarios and $4,438$ unseen-door scenarios in the simulator. Given the scarcity of prior research on door-opening tasks incorporating both tactile and visual information, we compare our method with the transformer-based multi-modal generator [17]. For fairness and practicality, we re-implement the latter without its auditory modality.
@@
Quantitative Results In order to accommodate the limitation of our real experiment, the robot used in our simulated experiment has a fixed base and is stationary. Therefore, the door is considered successfully opened if its angle only surpasses $30^{\circ}$. In addition, we establish $3$ levels of force thresholds (i.e. $\textbf{f}=5\text{N},~{}10\text{N}~{}\text{and}~{}15\text{N}$) to define SaR-95 and SaR-80 in order to evaluate the force safety performance of such involved states planning models more comprehensively. Tab. I presents the quantitative results of the models in both the seen-door and unseen-door scenarios discussed earlier. As shown, our method outperforms the others across nearly all metrics. This demonstrates that our method effectively ensures force safety during the robotic manipulation process and can generalize robustly to unseen scenarios.
Figure 3: Qualitative comparisons of different methods on three samples randomly selected from SafeDoorManip50k unseen-door scenarios with disturbance. Each sub-figure is annotated with two key indicators: the “blue line” represents the magnitude of harmful force applied by the robot’s end-effector, demonstrating the safety of the methods, while the “red line” shows the door opening angle, illustrating the effectiveness of the methods. The comparisons include: (a) Li et al. [17], (b) Our method without tactile calibration, and (c) Our method with tactile calibration.
@@
Q1: How does tactile calibration help safe state planning? As tactile calibration plays an essential role in our method, we conduct an ablation study to validate its importance by removing the force feedback input from our method. In the implementation, we directly bypass all operations associated with Eq. 5 during both the training and inference phases. As demonstrated in Tab. I, without tactile calibration, although our method still manages to successfully open doors, it fails to ensure force safety. More importantly, the absence of tactile calibration significantly impairs our method’s generalization capabilities, which indicates that vision-based state planning methods are inadequate for modeling the intricate dynamics inherent in robotic manipulation tasks, rendering them incapable of planning robustly in dynamic, unstructured environments.
TABLE II: Quantitative evaluation of different methods on our SafeDoorManip50k unseen-door scenarios with disturbance, highlighting the anti-disturbance capability of our method.
SuR (%) $\uparrow$ AHF (N) $\downarrow$ Threshold - 20 N
SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$
Li et al. [17] 3.83 17.92 0.00 7.83
Ours (V) 68.31 18.37 0.62 45.10
Ours (V+T) 95.18 9.59 28.25 89.25
[…]
V-C Real-world Experiments
@@ VI Conclusions
Acknowledgements
Special thanks to Prof. Yanding Wei (Zhejiang University) for real-world experiments support and helpful feedback on writing.

### v2

@@ 4 Dataset
5 Benchmark
5.1 Evaluation Metrics
@@
Success Rate (SuR) Unlike [li2024unidoormanip], we focus on manipulation force safety by defining SuR as the fraction of test scenarios in which the model completes the task without exceeding 20 N peak force—a threshold chosen under the assumption that forces above 20 N would mechanically damage both the robot and the manipulated object.
@@
Average Harmful Force (AHF) and Maximum Harmful Force (MHF) AHF and MHF is applied to evaluate the force-wise force safety of the state planning model in manipulation tasks. It is calculated as the average and maximum harmful force magnitude $\|\textbf{F}_{\texttt{harmful}}\|$ applied throughout each test process across all test scenarios respectively.
Safety Rate (SaR-95 and SaR-80) Safety Rates are utilized to evaluate the scenario-wise force safety of the state planning model in manipulation tasks. It is used to ensure that, most of the harmful force magnitudes during the operation remain relatively low, thereby protecting the robot and objects from being continuously exposed to high interaction forces. Since force safety of a state planning model only depends on the state generated by itself, rather than other states in that trajectory. Thus, we discretize the trajectories with the states planned along the way, i.e.
$\|\textbf{F}_{\texttt{harmful}}\|^{k}\leq\textbf{f},\quad\forall k\in[1,L]$ (6)
where $L$ denotes the length of states planned by the model and f denotes the force threshold. We evaluate the force safety of our state planning model using two metrics: SaR-95 and SaR-80. A test scenario is considered safe under the SaR-95 criterion if $\geq$95% of its generated states satisfies Eq. 6; similarly, it meets the SaR-80 criterion if $\geq$80% of the states satisfy this condition. Denoting by $\mathbf{Num}_{95\%\texttt{safe}}$ and $\mathbf{Num}_{80\%\texttt{safe}}$ the numbers of test scenarios meeting the SaR-95 and SaR-80 criteria respectively, and by $\mathbf{Num}_{\texttt{success}}$ the total number of successfully manipulated scenarios, the final metrics are defined as follows:
$\texttt{SaR-95}=\frac{\mathbf{Num}_{95\%\texttt{safe}}}{\mathbf{Num}_{\texttt{success}}},\;\;\texttt{SaR-80}=\frac{\mathbf{Num}_{80\%\texttt{safe}}}{\mathbf{Num}_{\texttt{success}}}$ (7)
5.2 Simulation Experiments
@@
Implementation Our proposed SafeDiff model is implemented based on the publicly available Decision Diffuser code base [ajay2022conditional]. The training and testing processes are conducted using an NVIDIA A100 Tensor Core GPU. We utilize the training demonstrations provided by our SafeDoorManip50k for safe state planning. The training configuration is as follows: batch size is $256$, total training epochs are $500$, an initial learning rate of $10^{-4}$ with a decay rate of $0.985$, and the application of an Exponential Moving Average (EMA) with a decay factor of $0.995$. During testing, we evaluate the performance of the safe state planning models under $4,580$ seen-door scenarios and $4,438$ unseen-door scenarios in the simulator. We compare our method with three representative works: the transformer-based multi-modal regression model [li2022see], the diffusion-based trajectory generator [li2024unidoormanip], and the action chunking transformer with tactile feedback [li2025hapticactbridginghumanintuition]. For fairness and practicality, we re-implement the latter without its auditory modality.
Table 1: Quantitative evaluation of our method and existing models on the simulation scenarios from our SafeDoorManip50k, highlighting the effectiveness of our method in safe state planning. Ours (V) denotes our method utilizing only visual data as input, while Ours (V+T) incorporates both visual data and tactile calibration. ✓and ✗ indicate whether the door manipulated is seen or unseen.
Seen (?) SuR (%) $\uparrow$ AHF (N) $\downarrow$ MHF (N) $\downarrow$ Threshold - 5 N Threshold - 10 N
SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$ SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$
Li et al. [li2022see] 69.50 7.68 19.05 0.10 0.66 6.09 43.12
Haptic-ACT [li2025hapticactbridginghumanintuition] 47.10 8.41 27.32 0.00 0.04 3.12 25.18
UniDoorManip [li2024unidoormanip] 49.50 9.78 22.10 0.00 0.04 0.53 7.00
Ours (V) 78.89 6.31 16.86 0.83 7.87 22.41 57.65
Ours (V+T) 80.07 5.07 15.03 6.10 25.28 49.25 78.73
Li et al. [li2022see] 68.09 7.51 18.80 0.00 0.66 8.82 47.55
Haptic-ACT [li2025hapticactbridginghumanintuition] 43.94 8.57 27.34 0.02 0.49 2.40 23.05
UniDoorManip [li2024unidoormanip] 52.70 9.47 21.65 0.00 0.00 0.75 11.82
Ours (V) 51.49 13.08 22.50 0.87 2.57 8.15 21.35
Ours (V+T) 81.03 5.08 14.59 5.13 24.90 55.54 79.33
@@
Quantitative Results In order to accommodate the limitation of our real experiment, the robot used in our simulated experiment has a fixed base and is stationary. Therefore, the door is considered successfully opened if its angle only surpasses $30^{\circ}$. In addition, we establish $2$ levels of force thresholds (i.e. $\textbf{f}=5\text{N}~\text{and}~15\text{N}$) to define SaR-95 and SaR-80 in order to evaluate the force safety performance of such involved states planning models more comprehensively. Tab. 1 presents the quantitative results of the models in both the seen-door and unseen-door scenarios discussed earlier. As shown, our method outperforms the others across nearly all metrics. This demonstrates that our method effectively ensures force safety during the robotic manipulation process and can generalize robustly to unseen scenarios.
@@
Q1: How does tactile calibration help safe state planning? As tactile calibration plays an essential role in our method, we conduct an ablation study to validate its importance by removing the force feedback input from our method. In the implementation, we directly bypass all operations associated with Eq. 5 during both the training and inference phases. As demonstrated in Tab. 1, without tactile calibration, although our method still manages to successfully open doors, it fails to ensure force safety. More importantly, the absence of tactile calibration significantly impairs our method’s generalization capabilities, which indicates that vision-based state planning methods are inadequate for modeling the intricate dynamics inherent in robotic manipulation tasks, rendering them incapable of planning robustly in dynamic, unstructured environments.
Q2: Does SafeDiff still work under environmental disturbances? The goal of the disturbance experiment is to observe whether the state planning methods can counteract the environmental disturbances, preventing their accumulation and ultimately avoiding failure in the robotic manipulation tasks. In the implementation, we tested the involved models using $4,438$ unseen-door scenarios from our SafeDoorManip50k dataset. And during the door-opening process, we applied a periodic impulsive ($1.5\text{Hz}$) disturbance with a positional deviation of $0.03$ meters. Some sample result is visualized in Fig. 5 of the appendix section C. The (a) is from Ours (V), which fails to overcome the disturbance, and (c) is from Ours (V+T). The (b), (d), (e) are from [li2022see], [li2025hapticactbridginghumanintuition], and [li2024unidoormanip] respectively. Our method with tactile calibration responds effectively to the disturbances, maintaining the harmful forces within a relatively small range, and ultimately succeeding in opening the door.
5.3 Real-world Experiments
@@ 6 Conclusions
Acknowledgments
This work was partially supported by Shenzhen Science and Technology Program (JCYJ20220818103001002), the Guangdong Key Laboratory of Big Data Analysis and Processing, Sun Yat-sen University, China, and by the High-performance Computing Public Platform (Shenzhen Campus) of Sun Yat-sen University.
The authors would like to express their sincere gratitude to Prof. Yanding Wei for supporting the real-world experiments and for providing valuable feedback on the manuscript.
Appendix A SafeDoorManip50k Dataset Details
A.1 Door assets production
The door body and handle assets are organized from [li2024unidoormanip]. See Fig. 4 for some door samples visualization. We employed the contact force interface within Isaac Gym to obtain the contact forces between the robot and the door. Notably, due to functional limitations of the Isaac Gym simulation engine, friction force (one source of the contact forces) cannot be read from the Isaac Gym’s API. Consequently, we had updated the collision mesh’s shape and parameter of the doorknob model, enabling accurate readings of the harmful forces in the horizontal direction.
Figure 4: Sample of simulation environments
A.2 Data collection configuration
Leveraging the parallel simulation capabilities of Isaac Gym, we simulate 10 distinct door environments per batch. The ground-truth labels for door-pulling trajectories are computed analytically, as the pose of the door handle in the world coordinate system can be directly queried via the Isaac Gym API. Given the handle pose, the target label at any specified door opening angle is determined by analytically calculating the handle’s position under the door opening angle with an angular offset.
To facilitate the model’s capability of implicit tactile calibration capabilities, we introduce temporally decaying random positional noise to the target coordinates of the robot’s end effector during execution. The noise direction is uniformly sampled from the interval $0,\;2\pi]$, while its magnitude is defined as $A=a\,e^{-k(t-t_{0})}$, where $a,e$ are randomized parameters, $t$ denotes the simulation timestep and $t_{0}$ marks the onset of the current noise perturbation cycle. We perform 5-6 perturbation cycles in each task.
A.2.1 Random strategies
We provide all random strategy configurations in this part. Here, $\texttt{Uniform}(a,\;b)$ denotes a uniform distribution over the range $[a,b]$, and $\texttt{Gaussian}(\mu,\;\sigma)$ denotes a Gaussian (normal) distribution with mean $\mu$ and standard deviation $\sigma$.
Table 2: Parameters and Sampling Distributions in the Simulation Environment
Parameter Sampling Distribution
Door type Uniformly sampled from the given door asset set
Door scale $\texttt{Uniform}(0.8,\;1.0)$
Door relative position offset (x, y, z) $\texttt{Gaussian}(0,\;0.08)\;m$, $\texttt{Gaussian}(0,\;0.06)\;m$,
$\texttt{Gaussian}(0,\;0.06)\;m$
Door hinge friction $\texttt{Uniform}(1,\;10)\;N$
Door hinge stiffness $\texttt{Uniform}(1,\;10)\;N/\mathrm{rad}$
Robot end-effector stiffness $\texttt{Gaussian}(3000,\;100)\;N/m$
Robot end-effector damping $\texttt{Gaussian}(100,\;10)\;N/(m/s)$
Environment light intensity $\texttt{Uniform}(0.3,\;0.7)$
Perturbation parameter $a$ $\texttt{Uniform}(3,\;18)\;mm$
Perturbation parameter $e$ $\texttt{Uniform}(1.2,\;1.8)$
Appendix B Real-world Experiment Details
B.1 Evaluation Settings
Following the simulation experiment, we establish two levels of force thresholds, $\textbf{F}_{\texttt{thres}}=10\text{N}~\text{and}~15\text{N}$, to define SaR-95 and SaR-80. We have omitted $\textbf{F}_{\texttt{thres}}=5\text{N}$ due to the real-world noise. In addition, a door is considered successfully opened if its angle exceeds $30^{\circ}$ while maximum force is smaller than 20N. Due to limited hardware resources, we were unable to conduct large-scale parallel real-world experiments; however, the results presented demonstrate our model’s effectiveness. During the experiment, three different types of door are used, where one for training and the other two for testing. Each test is repeated 10 times, with other experiment configurations (such as lighting and relative positions between robot and door) are randomized within a certain range. In addition, we chose Li et al. [li2022see] as the baseline for comparison because it achieved the best performance in simulation.
Table 3: Quantitative evaluation of our method and existing models on the real-world scenarios, highlighting the effectiveness of our safe states planning method in the real world. Ours (V+T) represents our method utilizing both visual data and tactile calibration as inputs. The symbols ✓and ✗ indicate whether the door manipulated is seen or if there is a disturbance present.
Seen (?) Disturbance (?) Training Set SuR (%) $\uparrow$ AHF (N) $\downarrow$ MHF (N) $\downarrow$ Threshold - 10 N Threshold - 15 N
SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$ SaR-95 (%) $\uparrow$ SaR-80 (%) $\uparrow$
Li et al. [li2022see] Real (100%) 100 9.038 21.781 0 0 100 100
Ours (V+T) Real (100%) 100 4.737 16.774 60 100 100 100
Ours (V+T) Sim (100%) + Real (20%) 100 3.763 18.581 100 100 100 100
Li et al. [li2022see] Real (100%) 100 10.786 23.880 0 0 60 90
Ours (V+T) Real (100%) 100 6.564 17.339 10 50 50 100
Ours (V+T) Sim (100%) + Real (20%) 100 4.709 17.338 10 60 60 100
Li et al. [li2022see] Real (100%) 100 18.803 31.855 0 0 0 0
Ours (V+T) Real (100%) 100 6.250 28.143 0 100 100 100
B.2 Quantitative Results
To evaluate the efficacy of the involved models in real-world robotic manipulation, we train them using the entire dataset of 110 real-world demonstrations (denoted as Real (100%)). Each model is then deployed and target points are sent to the robot for the door-opening process. The first and second rows of Tab. 3 show that our method ensures force safety more effectively in real-world robotic manipulation. We elaborate on other aspects in the following discussion.
Generalization. Similar as the simulation experiment, we attempted to open doors that it had not encountered during the fine-tuning stage. The 2nd and 5th rows of Tab. 3 demonstrate that our method more effectively in generalization performance. Upon further analysis of these experimental results, we observe that the model is more sensitive to changes in door size than to changes in door appearance. Specifically, the robot performs slightly worse on unseen door sizes compared to unseen door faces. However, our implicit tactile calibration successfully corrects the motion trajectory. This observation aligns with the bionic principles discussed in the introduction, further validating the effectiveness of our model design.
Anti-disturbance. While keeping the previously fine-tuned model unchanged, we manually introduce external disturbances during the door-opening process to assess the real-world disturbance resistance of the models involved. Rows 2 and 8 of Table 3 show that our method can effectively calibrate real-world disturbances online, compared to the baseline method [li2022see] at Row 7. This indirectly demonstrates that our model adeptly utilizes tactile information as gain-type negative feedback, continuously adjusting the vision-guided planned states. This capability enables robust adaptability to dynamic environmental changes, further validating the effectiveness of our model design.
Few-shot Sim-to-Real Transfer. To evaluate the efficacy of our model through few-shot fine-tuning, we initially train it using 700 sampled simulation demonstrations (denoted as Sim). We then fine-tune it using only 20% of the 110 real-world demonstrations (denoted as Real (20%)), before deploying on the robot to guide the door-opening process. Rows 3 and 6 of Table 3 indicate that our few-shot fine-tuned model outperforms the model trained exclusively on the 110 real-world demonstrations. This superior performance can be attributed to the simulated environment’s ability to provide a more intricate and diverse array of training demonstrations, underscoring the importance of our large-scale multimodal simulation dataset, SafeDoorManip50k. This confirms the substantial contribution of such a dataset in enhancing model robustness and adaptability.
Appendix C Supplementary Figures
Figure 5: Quantitative evaluation of different methods on our SafeDoorManip50k unseen-door scenarios with disturbance, highlighting the anti-disturbance capability of our method.